# Just-In-Time Objectives: A General Approach for Specialized AI Interactions

### v1

@@ 3.1.2. Evaluate: Applying JIT Objectives to Optimize Evaluation
To try to get better feedback, Mary tries a standard LLM-as-judge setup to evaluate feedback options. She creates a rubric that assesses the overall quality and intellectual rigor of the feedback to select for higher-level feedback on the content and ideas rather than presentation. Without a just-in-time objective, she finds that many of the feedback options receive similarly high scores, and do not differentiate themselves.
The eval_objective operator applies a just-in-time objective to a provided evaluator to optimize its assessment towards a user’s goals. This can again be implemented simply by adding the induced objective JSON specification (Figure 3) to an existing evaluation prompt.
Adding the JIT objective to her LLM-as-judge prompt, Mary gets a larger spread of scores. For example, scoring feedback on the induced JIT objective of “Strengthen the narrative argument” allowed the system to filter out feedback that merely mentioned the importance of narratives, but did not provide concrete strategies on how to incorporate a narrative into her draft.
Added on to a recommendation module that selects the highest-relevance candidates, eval_objective can steer the system towards goal-aligned options. Added on to a critique module that produces feedback on generated candidates, eval_objective can aid iterative refinement loops by producing goal-based critiques that feed back into a generator to improve its next round of candidates. Added on to a sampling module that selects the best of N generations, eval_objective can serve as a verifier for test-time scaling towards user goals.
Figure 4. Inducing Just-In-Time Objectives. The core architecture depends on a single upstream step of objective induction that infers user goal(s) based on observed context and instantiates each goal as a just-in-time objective, which can later be applied to optimize generation (gen_objective) and evaluation (eval_objective). :
Figure 5. Poppins incorporates just-in-time objectives to power UI generation. The Poppins system instantiates our just-in-time objectives architecture by inducing objectives from a user’s context (e.g., their screen or text document). Then, generation and evaluation steps steer objective application for Tool Generation (Poppins-tools) and Expert Generation (Poppins-experts) to automatically generate a tailored output for the user’s task (interactive tools or expert responses, respectively).:
@@ 3.1.3. Inducing JIT Objectives
We’ve described how gen_objective and eval_objective apply a JIT objective to generation and evaluation tasks—but how do we produce this objective in the first place? The objective induction step aims to infer likely user goals and translate them into just-in-time objectives. We define a goal as a user’s intention or aim for a window of time. Meanwhile, we define an objective as a system’s operationalization of the user goal into a form that can be used to guide optimization.
@@
We perform objective induction by taking in user input data in the form of text, images, or file attachments. This input user context may take the form of a text string (e.g., content from Overleaf or Google Docs), a workspace screenshot (e.g., the user’s full browser window with a site like Figma, Google Slides, or Google Sheets, or their full desktop with an application like VSCode or Microsoft Word), or an attached file (e.g., an image of a poster draft, a PDF of a research article). We then prompt the LLM to induce objectives: we include the input user context, ask the system to infer goals, and request output in the form of the objective JSON specification. Our objective induction prompt includes a chain-of-thought process that has the model reflect on factors including the user’s task domain, stage of work completion, potential audience, ideal final task output, and anticipated reaction to assistance (full prompts in Appendix A).
@@ 3.2. Poppins: UI Generation Powered by Just-In-Time Objectives
Having outlined how we construct just-in-time objectives, we demonstrate how system developers can apply the just-in-time objectives architecture to steer LLMs in interactive systems. We develop Poppins, a system for UI generation that incorporates just-in-time objectives. Poppins is a browser extension and web application that observes user screens and, when enabled, automatically generates outputs to assist the user in their specific task. [^2] Much like our namesake Mary Poppins, whose bag always carried the right tool for the particular task at hand. Critically, rather than requiring users to devote time and cognitive effort to craft requests for customized assistance, our system recognizes user needs and produces outputs that users might not have time to create, or might not have thought to create. Our system can take on this alternate task formulation only because it builds on just-in-time objectives that set automated generation and evaluation processes on the right user-aligned course. Poppins supports two types of generated assistance: (1) Expertise generation with Poppins-experts and (2) Tool generation with Poppins-tools. These two types of assistance demonstrate how LLMs can better support common workflows with JIT objectives: (1) feedback and suggestion on user input, and (2) on-demand interface construction and customization.
Figure 6. Poppins-experts uses just-in-time objectives to generate relevant model expertise, select an appropriate output format (e.g., Feedback, Brainstorm), and generate model responses drawing on the specified expertise.:
@@ 3.2.1. Input Pipeline
Mary is working on a CHI paper draft in Overleaf, specifically the System section. She’s been looking at it so long that she can’t tell if it makes sense. Could Poppins help? She opens the extension sidebar and clicks the button to capture her window. Under the hood, based on the text being edited, the system infers candidate goals of “Enhance technical clarity” (weight: 9) and “Strengthen evaluation presentation” (weight: 8) and selects the higher-weighted first option to translate into a just-in-time objective.
While these input modes provide a lightweight, non-invasive entrypoint into users workflows, our system is flexible to support other input sources, so it could be extended to support streams of periodic screen captures, continuous video recordings, webpage interaction logs, or revision histories for more rich and domain-specific context to inform objective induction.
@@ 3.2.2. Expertise Generation with Poppins-experts
The first form of assistance that Poppins implements is expertise generation: it leverages the JIT objective to identify experts, perspectives, and areas of expertise that might be most helpful for assistance.
@@
Using the JIT objective, Poppins’s expertise generator produces an expert specification that includes a name, description, and several paragraphs of detailed background material retrieved using LLM search, such as relevant publications, talks, and projects; specific expertise areas and methodologies; and key ideas or quotes to consider. Poppins appends a gen_objective operator to steer expertise generation towards the user’s goals. Then, it performs evaluation with an eval_objective operator to score the candidate experts and select the top goal-aligned suggestions.
@@
Mary has Poppins-experts enabled, so the system applies the just-in-time objective to the expertise generator to surface background knowledge relevant to her paper and goal of “enhancing technical clarity.” This produces a “Technical Writing Specialist,” “Systems Architecture Expert,” and “Human-AI Interaction Researcher,” each of which is supported by detailed background information referencing particular academic papers, researchers, and concepts that bridge the expertise area and the content of Mary’s paper (Figure 6).
Figure 7. Poppins-tools applies just-in-time objectives to generate tool designs, optionally generate relevant model expertise, and generate UI code to instantiate the tool idea.:
@@
After model expertise generation, Poppins-experts uses an output evaluator to determine the appropriate output format for this expert. While we could leave out this step and simply prompt the model using the expert specification, setting the desired output format affords greater control over expert output. We implement an initial set of output formats, each of which steers prompts to the experts and is mapped to a UI template to render the expert outputs: Feedback (generates feedback and a synthesis of common themes and unique perspectives), Brainstorm (generates brainstorming ideas based on input content), and Line Editor (interactively generates line-level edits on user-highlighted text). The output evaluator takes in the user context and selects the most relevant output format; Poppins adds on an eval_objective operator to account for user goals in this selection.
@@
Based on Mary’s Overleaf context and induced objective, Poppins-experts selects the Feedback output format and loads the associated Feedback UI in the extension sidebar with her paper content pre-loaded as input. Mary views the generated experts and clicks the “Run“ button to proceed with gathering feedback. The resulting output surfaces helpful feedback from the Human-AI Interaction Researcher that the section “could be clearer about the specific metrics used to assess objective accuracy and usefulness,” and that the paper would benefit from “deeper discussion of the cognitive load implications” of users reviewing system’s automatic inferences. Meanwhile, the System Architecture Expert raises that “internal component interfaces are well-defined,” but the paper “could better specify how external systems would integrate with this architecture,” which Mary flags as a potential topic in the Discussion.
@@ 3.2.3. Tool Generation with Poppins-tools
Poppins demonstrates a second form of assistance via tool generation: based again on a user’s context, Poppins-tools attempts to design and synthesize a code implementation for an interactive tool to assist the user.
@@
After tool design generation, Poppins-tools runs a UI code generator that takes in the tool specification and optional expert specifications and generates a code implementation of the tool idea. This generator formulates a request for a standalone web interface implemented as a Svelte component (see prompts in Appendix A), and Poppins adds a gen_objective operator to encourage goal-alignment. To provide more reliable model behavior around LLM integration, we provide the system with access to a library of LLM helper functions (with functions like getExperts, promptExpert, promptGeneral). These allow the system to more easily formulate LLM calls, especially to consistently incorporate the requested expertise if useful to the induced tool. Given the high cost of generating candidates for the UI generation task using more advanced models, the UI code evaluator is a critique module that provides feedback to refine the current candidate rather than selecting over a pool of candidates. The evaluator performs checks for usability, design quality, and bug-free code, and Poppins adds an eval_objective operator to vet goal alignment. This step produces a final code implementation, which we render in Poppins.
@@ 3.3. System Limitations
We note several limitations of the just-in-time objectives architecture and Poppins system. First, a just-in-time approach necessarily incurs a time cost since users must wait for the objectives to be induced and applied, which currently takes 1-3 minutes. Since this is much slower than a standard LLM chat response, JIT objectives may be appropriate only when the time investment is worthwhile for a higher quality response. Our method is heavily reliant on the model to accurately infer and apply user objectives. While our evaluation finds that objectives are highly accurate, LLMs may have inductive biases such that they are less likely to induce certain objectives, or less performant when applying certain objectives, warranting further evaluation before widespread release. The architecture has minimal visibility into a user via an individual snapshot of context, which can result in some inaccurate induced objectives. While we can gather more data, there are limits to this strategy: there is no way to instrument a user to capture their full life context, and gathering more data can dramatically increase privacy and security risks for users.
@@ 4. EVALUATION: Assessing the Accuracy and Utility of Just-In-Time Objectives
Our evaluation has two main aims: (RQ1) to determine whether our just-in-time objectives are accurate and useful, and (RQ2) to investigate whether systems that incorporate just-in-time objectives are relevant and useful to users’ needs. To assess RQ1, we conduct experiments where participants evaluate just-in-time objectives induced from submitted screenshots of their own browsing behavior (Section 4). We assess RQ2 by conducting 1-hour in-lab sessions where participants receive assistance on their own writing task with a baseline LLM and the Poppins system that incorporates just-in-time objectives (Section 5).
@@ 4.1.1. Study 1: Assessing Accuracy & Utility
For the first study, we gather screenshots from participants’ daily tasks as input contexts to induce just-in-time objectives. We recruit participants from university mailing lists and social media in accordance with our institution’s IRB. Participants first submit a pre-survey with background information on their LLM use, along with a link to a document they will use for their writing task during the in-lab study session (Section 5). In total, 17 participants took part in our study (see Appendix B.3 for participant background information). Participants were compensated with a \$35 Amazon gift card for completing the hour-long lab study and an additional \$15 for participating in an accompanying screenshot study.
We asked participants to complete a screenshot data collection task that is used for an optional study, and 14 of 17 participants completed this additional study. To gather screenshot data, participants install a lightweight Chrome extension that allows them to capture their browser (a screenshot of the visible window and extracted text content from the page) and submit the data to our study. We instructed participants to share any browser content that was representative of their daily tasks and work. All surveys and instructions are included in Appendix B.1. We collect browser data over the course of three days, resulting in an average of 11 screenshots per participant. The submitted browser data spanned a range of interfaces (e.g., Google Docs, Google Slides, LLM chats, and academic paper databases), task types (e.g., presentation feedback, writing feedback, text comprehension, brainstorming), and topic areas (e.g., chemistry, neuroscience, natural language processing, linguistics, design, energy policy, TA training).
@@
Figure 8. Study 1: Induced from participants’ own input contexts (70 inputs from N=14N={14}{} participants), JIT objectives are accurate, useful, and produce outputs preferred to a baseline LLM. Participants rate the induced objective as accurate and useful in the vast majority of cases (A, B). Experts, tool designs, and feedback generated with a JIT objective are preferred over that of a baseline LLM in 71–86% of cases (C), and sampling JIT generations with a JIT evaluator produces more preferred outputs (D).:
@@
For each participant, we run a data selection script on their submitted browser data to select five items as inputs to induce just-in-time objectives, providing a total of $n=70$ input contexts across participants. The script enforces criteria about (1) suitability: whether the provided item would benefit from AI assistance, (2) quality: whether the screenshot has enough context to understand the user’s task, and (3) divergence: whether the item sufficiently diverges from other suitable items, to encourage a spread of differing tasks for each user (e.g., not all academic writing tasks). Thus, the script allows us to focus our study session around a set of informative real-life “moments” at which our system might intervene, while providing some consistency across participants. The selected images are used in a 20-minute study with two sections:
@@
  • Objectives and generator task. The first section allows us to assess the quality of our just-in-time objectives and JIT generator. For four of the five selected inputs, participants assess the usefulness and accuracy of the highest-weighted induced objective based on the input. Then, they provide pairwise preferences for experts, tool designs, and feedback generated using a just-in-time objective versus a baseline LLM (Claude Sonnet 3.7).
@@
  • Evaluator best-of-N task. The second section allows us to investigate the efficacy of our JIT evaluator. A common practice for producing high quality output from an LLM involves sampling many ($N$) candidate outputs and selecting the best (best-of-NN(Stiennon et al., 2020; Snell et al., 2024; Beirami et al., 2024). A strong JIT evaluator should produce better outputs as $N$ increases—the intuition here is that a model gets more “attempts” to produce something strong, with the JIT evaluator returning the best final candidate. To evaluate this, participants provide pairwise preferences on output evaluated with a just-in-time objective, with progressively increasing sample sizes. For each sample size ($N=1,10,100$), we generate $N$ candidates using a JIT generator, and then select a single best candidate using a JIT evaluator. The best candidate from each sample is presented to the user (the “best-of-N” output).
@@
Figure 9. Study 2: Across an expanded set of participant-submitted input contexts (410 inputs from N=205N={205}{} participants), just-in-time objectives are accurate, useful, and produce outputs preferred to a baseline LLM. Participants rate the induced objective as accurate and useful in the vast majority of cases (A, B). Experts, tool designs, and feedback generated with a just-in-time objective are preferred over that of a baseline LLM (C), and sampling over JIT generations with a JIT evaluator produces more preferred outputs, though with marginal benefit beyond 10 samples (D).:

### v2

@@ 3.1.2. Evaluate: Applying JIT Objectives to Optimize Evaluation
To try to get better feedback, Mary tries a standard LLM-as-a-judge setup to evaluate feedback options. She creates a rubric that assesses the overall quality and intellectual rigor of the feedback to select for higher-level feedback on the content and ideas rather than presentation. Without a just-in-time objective, she finds that many of the feedback options receive similarly high scores, and do not differentiate themselves.
The eval_objective operator applies a JIT objective to a provided evaluator to optimize its assessment towards a user’s goals. This can again be implemented simply by adding the induced objective JSON specification (Figure 3) to an existing evaluation prompt.
Adding the JIT objective to her LLM-as-a-judge prompt, Mary gets a larger spread of scores. For example, scoring feedback on the induced JIT objective of “Strengthen the narrative argument” allowed the system to filter out feedback that merely mentioned the importance of narratives, but did not provide concrete strategies on how to incorporate a narrative into her draft.
Added on to a recommendation module that selects the highest-relevance candidates, eval_objective can steer the system towards objective-aligned options. Added on to a critique module that produces feedback on generated candidates, eval_objective can aid iterative refinement loops by producing objective-based critiques that feed back into a generator to improve its next round of candidates. Added on to a sampling module that selects the best of N generations, eval_objective can serve as a verifier for test-time scaling towards user goals.
Figure 4. Inducing Just-In-Time Objectives. The core architecture depends on a single upstream step of objective induction that infers user goal(s) based on observed context and instantiates each goal as a just-in-time objective, which can later be applied to optimize generation (gen_objective) and evaluation (eval_objective).
Architecture diagram showing how user context observation leads to an objective, which produces generation and evaluation operators that can steer downstream AI systems.
@@ 3.1.3. Inducing JIT Objectives
We’ve described how gen_objective and eval_objective apply a JIT objective to generation and evaluation tasks, but how do we produce this objective in the first place? The objective induction step aims to infer likely user goals and translate them into just-in-time objectives. We define a goal as a user’s intention or aim for a window of time. Meanwhile, we define an objective as a system’s operationalization of the user goal into a form that can be used to guide optimization.
@@
We perform objective induction by taking in user input data in the form of text, images, or file attachments. This input user context may take the form of a text string (e.g., content from Overleaf or Google Docs), a workspace screenshot (e.g., the user’s full browser window with a site like Figma, Google Slides, or Google Sheets, or their full desktop with an application like VSCode or Microsoft Word), or an attached file (e.g., an image of a poster draft, a PDF of a research article). We then prompt a vision language model (VLM) to induce objectives: we include the input user context, ask the system to infer goals, and request output in the form of the objective JSON specification (Figure 3). Our objective induction prompt includes a chain-of-thought process that has the model reflect on factors including the user’s task domain, stage of work completion, potential audience, ideal final task output, and anticipated reaction to assistance (full prompts in Appendix A).
@@
Figure 5. Poppins incorporates just-in-time objectives to power UI generation. The Poppins system instantiates our just-in-time objectives architecture by inducing objectives from a user’s context (e.g., their screen or text document). Then, generation and evaluation steps steer objective application for Tool Generation (Poppins-tools) and Expert Generation (Poppins-experts) to automatically generate a tailored output for the user’s task (interactive software tools or expert responses, respectively).
A system diagram showing how Poppins uses just-in-time objectives to power two main functions: Poppins-tools (interactive tool generation) and Poppins-experts (expert response generation). The architecture shows user context input flowing through objective induction, then through generation and evaluation steps guided by the objectives, ultimately producing tailored outputs for the user’s specific task.
@@ 3.2. Poppins: UI Generation Powered by Just-In-Time Objectives
Having outlined how we construct just-in-time objectives, we demonstrate how system developers can apply the just-in-time objectives architecture to steer LLMs in interactive systems. We develop Poppins, a system for UI generation that incorporates just-in-time objectives (Figure 5). Poppins is a browser extension and web application that observes user screens and, when enabled, automatically generates outputs to assist the user in their specific task. [^3] Much like our namesake Mary Poppins, whose bag always carried the right tool for the particular task at hand. Critically, rather than requiring users to devote time and cognitive effort to craft requests for customized assistance, our system recognizes user needs and produces outputs that users might not have time to create, or might not have thought to create. Our system can take on this alternate task formulation only because it builds on just-in-time objectives that set automated generation and evaluation processes on the right user-aligned course. To demonstrate how JIT objectives can integrate into existing LLM systems of interest to the HCI community, Poppins supports two types of generated assistance: (1) Expertise generation with Poppins-experts and (2) Tool generation with Poppins-tools. These two types of assistance demonstrate how LLMs can better support common workflows with JIT objectives: (1) feedback on user input with expert perspectives (e.g., Benharrak et al., 2024; D’Arcy et al., 2024; Choi et al., 2025), and (2) on-demand interface construction and customization (e.g., Cheng et al., 2024; Vaithilingam et al., 2024; Chen et al., 2025; Cao et al., 2025).
Figure 6. Poppins-experts uses just-in-time objectives to generate relevant model expertise, select an appropriate output format (e.g., Feedback, Brainstorm), and generate model responses drawing on the specified expertise. Users can view and modify all objectives, output formats, and experts.
Poppins-experts interface displaying generated expert profiles and feedback outputs. A screenshot of the browser extension interface showing three generated expert profiles: Technical Writing Specialist, Systems Architecture Expert, and Human-AI Interaction Researcher. The interface shows the objectives, output format, and experts selected to create this output. The output interface shows input fields for the user’s content and text responses from each expert, along with a synthesis of main commonalities and differences among the experts.
@@ 3.2.1. Input Pipeline
Mary is working on a CHI paper draft in Overleaf, specifically the System section. She’s been looking at it so long that she can’t tell if it makes sense. Could Poppins help? She opens the extension sidebar and clicks the button to capture her window. Under the hood, based on the text being edited, the system infers candidate objectives of “Enhance technical clarity” (weight: 9) and “Strengthen evaluation presentation” (weight: 8) and selects the higher-weighted first option to translate into a just-in-time objective.
While these input modes provide a lightweight, non-invasive entrypoint into user workflows, our system is flexible to support other input sources, so it could be extended to support streams of periodic screen captures, continuous video recordings, webpage interaction logs, or revision histories for more rich and domain-specific context to inform objective induction.
@@ 3.2.2. Expertise Generation with Poppins-experts
The first form of assistance that Poppins implements is expertise generation: it leverages the JIT objective to identify experts, perspectives, and areas of expertise that might be most helpful for assistance (Figure 6).
@@
Using the JIT objective, Poppins’s expertise generator produces an expert specification that includes a name, description, and several paragraphs of detailed background material retrieved using LLM search, such as relevant publications, talks, and projects; specific expertise areas and methodologies; and key ideas or quotes to consider. Poppins appends a gen_objective operator to steer expertise generation towards the user’s objective. Then, it performs evaluation with an eval_objective operator to score the candidate experts and select the top objective-aligned suggestions.
@@
Mary has Poppins-experts enabled, so the system applies the just-in-time objective to the expertise generator to surface background knowledge relevant to her paper and objective of “enhancing technical clarity.” This produces a “Technical Writing Specialist,” “Systems Architecture Expert,” and “Human-AI Interaction Researcher,” each of which is supported by detailed background information referencing particular academic papers, researchers, and concepts that bridge the expertise area and the content of Mary’s paper (Figure 6).
@@
After model expertise generation, Poppins-experts uses an output evaluator to determine the appropriate output format for this expert. While we could leave out this step and simply prompt the model using the expert specification, setting the desired output format affords greater control over expert output. We implement an initial set of output formats, each of which steers prompts to the experts and is mapped to a UI template to render the expert outputs: Feedback (generates feedback and a synthesis of common themes and unique perspectives), Brainstorm (generates brainstorming ideas based on input content), and Line Editor (interactively generates line-level edits on user-highlighted text). The output evaluator takes in the user context and selects the most relevant output format; Poppins adds on an eval_objective operator to account for the user objective in this selection.
@@
Based on Mary’s Overleaf context and induced objective, Poppins-experts selects the Feedback output format and loads the associated Feedback UI in the extension sidebar with her paper content pre-loaded as input. Mary views the generated experts and clicks the “Run” button to proceed with gathering feedback. The resulting output surfaces helpful feedback from the Human-AI Interaction Researcher that the section “could be clearer about the specific metrics used to assess objective accuracy and usefulness,” and that the paper would benefit from “deeper discussion of the cognitive load implications” of users reviewing system’s automatic inferences. Meanwhile, the System Architecture Expert raises that “internal component interfaces are well-defined,” but the paper “could better specify how external systems would integrate with this architecture,” which Mary flags as a potential topic in the Discussion.
@@
By default, Poppins-experts presents an automated workflow to both generate experts and expert response outputs. However, users can optionally modify any of the intermediate generations (Objectives, Experts, and Output formats) with several actions (Figure 6). They can Select other options that did not receive the highest score, Edit existing options by rewriting system-generated content in the form fields (e.g., expert name, description, background material), Add new options and manually author all form fields, and Delete any unwanted options.
Figure 7. Poppins-tools applies just-in-time objectives to generate tool designs, optionally generate relevant model expertise, and generate UI code to instantiate the tool idea. Users can view and modify all objectives, tool designs, and (optional) experts. Here, the system induced an objective of “Create clear visual representations of the AI system.” Poppins-tools then generated a Component Relationship Diagram Builder, a drag-and-drop diagramming UI to visualize system components, experiment with layouts, and request expert feedback on the system architecture diagram.
Poppins-tools interface showing generated tool specifications and implementation. A screenshot displaying a Component Relationship Diagram Builder tool generated by the system. The interface shows the objectives, tools, and experts selected to create this tool. The implemented tool appears as a diagram canvas with draggable component blocks and connection options, along with an AI Architecture Insights button for expert feedback.
@@ 3.2.3. Tool Generation with Poppins-tools
Poppins demonstrates a second form of assistance via tool generation: based again on a user’s context, Poppins-tools attempts to design and synthesize a code implementation for an interactive tool to assist the user (Figure 7). Here, we define a “tool” as interactive software that assists the user, as opposed to static text responses (e.g., chat reply, feedback, draft document) that are typical of LLM chat systems and encapsulated by Poppins-experts. To demonstrate the space of Poppins-tools outputs, we include examples (Figures 11, 15, 16).
@@
After tool design generation, Poppins-tools runs a UI code generator that takes in the tool specification and optional expert specifications and generates a code implementation of the tool idea. This generator formulates a request for a standalone web interface implemented as a Svelte component (see prompts in Appendix A), and Poppins adds a gen_objective operator to encourage objective alignment. To provide more reliable model behavior around LLM integration, we provide the system with access to a library of LLM helper functions (with functions like getExperts, promptExpert, promptGeneral). These allow the system to more easily formulate LLM calls, especially to consistently incorporate the requested expertise if useful to the induced tool. Given the high cost of generating candidates for the UI generation task using more advanced models, the UI code evaluator is a critique module that provides feedback to refine the current candidate rather than selecting over a pool of candidates. The evaluator performs checks for usability, design quality, and bug-free code, and Poppins adds an eval_objective operator to vet objective alignment. This step produces a final code implementation, which we render in Poppins.
@@
Similar to Poppins-experts, Poppins-tools by default automatically generates tool designs and the tool UI code. Users can optionally modify intermediate generations (Objectives, Tools, and Experts) with the same actions: Select other options, Edit an existing option (e.g., rewriting the tool input type, output type, description), Add a new option, or Delete an unwanted option (Figure 7).
@@ 3.3. System Limitations
We note several limitations of the just-in-time objectives architecture and Poppins system. First, a just-in-time approach necessarily incurs a time cost since users must wait for the objectives to be induced and applied, which currently takes 1-3 minutes. Since this is much slower than a standard LLM chat response, JIT objectives may be appropriate only when the time investment is worthwhile for a higher quality response. Our method is heavily reliant on the model to accurately infer and apply user objectives. While our evaluation finds that objectives are highly accurate, LLMs may have inductive biases such that they are less likely to induce certain objectives, or less performant when applying certain objectives, warranting further evaluation before widespread release.
The architecture has minimal visibility into a user via an individual snapshot of context, which can result in some inaccurate induced objectives. While this lightweight approach demonstrates a “floor” of performance, a single screenshot has limited visibility into complex tasks that require more context over time, or subtle needs that require more observations to understand. While we can gather more data, there are limits to this strategy: there is no way to instrument a user to capture their full life context, and gathering more data can dramatically increase privacy and security risks for users. Longer context windows increase coverage of the necessary information to accurately infer objectives, but also introduce new challenges of effectively filtering out noise and handling contradictory evidence.
@@ 4. EVALUATION: Assessing the Accuracy and Utility of Just-In-Time Objectives
Our evaluation has two main aims: (RQ1) to determine whether our just-in-time objectives are accurate and useful, and (RQ2) to investigate whether systems that incorporate just-in-time objectives are relevant and useful to users’ needs. To assess RQ1, we conduct experiments where participants evaluate just-in-time objectives induced from submitted screenshots of their own browsing behavior (Section 4). We assess RQ2 by conducting 1-hour in-lab sessions where participants receive assistance on their own writing task with a baseline LLM and the Poppins system that incorporates just-in-time objectives (Section 5). To establish a “floor” of performance, all evaluations used a single screenshot, the minimal context window, for objective induction. This evaluation setup allows us to assess JIT objective performance on a minimally-invasive setting of individual user screens.
@@ 4.1.1. Study 1: Assessing Accuracy & Utility
For the first study, we gather screenshots from participants’ daily tasks as input contexts to induce just-in-time objectives. We recruit participants from university mailing lists and social media in accordance with our institution’s IRB. Participants first submit a pre-survey with background information on their LLM use, along with a link to a document they will use for their writing task during the in-lab study session (Section 5). In total, 17 participants took part in our study (see Appendix C.3 for participant background information). Participants were compensated with a \$35 Amazon gift card for completing the hour-long lab study and an additional \$15 for participating in an accompanying screenshot study.
@@
Figure 8. Study 1: Induced from participants’ own input contexts (70 inputs from N=14N={14}{} participants), just-in-time objectives are accurate, useful, and produce outputs preferred to a baseline LLM. The vast majority of participants rate the induced objective as accurate and useful (A, B). Experts, tool designs, and feedback generated with a JIT objective are preferred over that of a baseline LLM in 71–86% of contexts (C), and sampling over JIT generations with a JIT evaluator produces preferred outputs (D).
Survey results from 14 participants showing just-in-time objectives are accurate, useful, and produce preferred outputs. Four bar charts showing evaluation results: (A) Objective accuracy ratings on a 7-point Likert scale with 75% rated as accurate or very accurate, (B) Objective usefulness ratings with 75% rated as useful or very useful, (C) Win rates for just-in-time objectives versus baseline large language model across experts (86%), tool designs (73%), and feedback (71%), and (D) Best-of-N evaluation showing increasing preference for higher sample sizes with just-in-time evaluator.
We asked participants to complete a screenshot data collection task that is used for an optional study, and 14 of 17 participants completed this additional study. To gather screenshot data, participants install a lightweight Chrome extension that allows them to capture their browser (a screenshot of the visible window and extracted text content from the page) and submit the data to our study. We instructed participants to share any browser content that was representative of their daily tasks and work. All surveys and instructions are included in Appendix C.1. We collect browser data over the course of three days, resulting in an average of 11 screenshots per participant. The submitted browser data spanned a range of interfaces (e.g., Google Docs, Google Slides, LLM chats, and academic paper databases), task types (e.g., presentation feedback, writing feedback, text comprehension, brainstorming), and topic areas (e.g., chemistry, neuroscience, natural language processing, linguistics, design, energy policy, TA training).
@@
For each participant, we run a data selection script on their submitted browser data to select five items as inputs to induce just-in-time objectives, providing a total of $n=70$ input contexts across participants. The script enforces criteria about (1) suitability: whether the provided item would benefit from AI assistance, (2) quality: whether the screenshot has enough context to understand the user’s task, and (3) divergence: whether the item sufficiently diverges from other suitable items, to encourage a spread of differing tasks for each user (e.g., so that a user’s tasks are not all academic writing). Thus, the script allows us to focus our study session around a set of informative real-life moments at which our system might intervene, while providing some consistency across participants. The selected images are used to generate a 20-minute study with two sections:
@@
  • Objectives and generator task. The first section allows us to assess the quality of our just-in-time objectives and JIT generator. For four of the five selected inputs, participants assess the usefulness and accuracy of the highest-weighted induced objective based on the input. Then, they provide pairwise preferences for experts, tool designs, and feedback generated using a just-in-time objective versus a baseline. This baseline condition received the same user screenshot, prompt, and LLM (Claude Sonnet 3.7) as the JIT objective condition, but just did not include an induced objective.
@@
  • Evaluator best-of-N task. The second section allows us to investigate the efficacy of our JIT evaluator added on to a JIT generator. A common practice for producing high quality output from an LLM involves sampling many ($N$) candidate outputs and selecting the best (best-of-NN(Stiennon et al., 2020; Snell et al., 2024; Beirami et al., 2024). A strong JIT evaluator should produce better outputs as $N$ increases. The intuition here is that a model gets more “attempts” to produce something strong, with the JIT evaluator returning the best final candidate. To evaluate this, participants provide pairwise preferences on output evaluated with a just-in-time objective, with progressively increasing sample sizes. For each sample size ($N=1,10,100$), we generate $N$ candidates using a JIT generator, and then select a single best candidate using a JIT evaluator. The best candidate from each sample is presented to the user (the “best-of-N” output).
@@
Figure 9. Study 2: Across an expanded set of participant-submitted input contexts (410 inputs from N=205N={205}{} participants), JIT objectives are accurate, useful, and produce outputs preferred to a baseline LLM. The vast majority of participants rate the induced objective as accurate and useful (A, B). Experts, tool designs, and feedback generated with a just-in-time objective are preferred over that of a baseline LLM (C), and sampling over JIT generations with a JIT evaluator produces more preferred outputs, though with marginal benefit beyond 10 samples (D).
Expanded survey results from 205 participants confirming just-in-time objective effectiveness across broader contexts. Four bar charts mirroring Figure 8 structure but with larger sample size: (A) 77% of objectives rated accurate or very accurate, (B) 80% rated useful or very useful, (C) Win rates of 71%, 66%, and 66% for experts, tool designs, and feedback respectively, and (D) Best-of-N results showing diminishing returns beyond 10 samples.